OpenStudy (riotschon):

An equation is shown below: 5(2x - 8) + 15 = -15 Write the steps you will use to solve the equation and explain each step.

OpenStudy (anonymous):

First distribute the 5 and simplify then combine like x^0 terms and solve for x by division.

alones (alones):

Step 1: Simplify both sides of the equation. 5(2x−8)+15=−15 Simplify: (Show steps) 10x−25=−15 Step 2: Add 25 to both sides. 10x−25+25=−15+25 10x=10 Step 3: Divide both sides by 10. 10x/10=10/10 x=1

OpenStudy (mathmale):

@AloneS, The purpose of this problem was to challenge the student to analyze the steps necessary to solve this equation and then follow those steps to finding it. Doing this work for that person has little positive value, unless you involve that person.